# Diffusion-Based mmWave Radar Point Cloud Enhancement Driven by Range Images

### v1

@@ Abstract
Millimeter-wave (mmWave) radar has attracted significant attention in robotics and autonomous driving. However, despite the perception stability in harsh environments, the point cloud generated by mmWave radar is relatively sparse while containing significant noise, which limits its further development. Traditional mmWave radar enhancement approaches often struggle to leverage the effectiveness of diffusion models in super-resolution, largely due to the unnatural range-azimuth heatmap (RAH) or bird’s eye view (BEV) representation. To overcome this limitation, we propose a novel method that pioneers the application of fusing range images with image diffusion models, achieving accurate and dense mmWave radar point clouds that are similar to LiDAR. Benefitting from the projection that aligns with human observation, the range image representation of mmWave radar is close to natural images, allowing the knowledge from pre-trained image diffusion models to be effectively transferred, significantly improving the overall performance. Extensive evaluations on both public datasets and self-constructed datasets demonstrate that our approach provides substantial improvements, establishing a new state-of-the-art performance in generating truly three-dimensional LiDAR-like point clouds via mmWave radar.
@@ I Introduction
Figure 1: The comparative demonstration includes LiDAR range images and point clouds, as well as the range images and high-quality millimeter-wave radar point clouds obtained through the proposed method. The red and green dashed boxes highlight the flowerbed and bench in the current scene, respectively. We can easily identify object categories and obtain their geometric structures through range images, demonstrating that range images preserve rich semantic and geometric information. As a proxy representation of point clouds, it enables the knowledge of pre-trained image diffusion models to be fully leveraged in point cloud super-resolution tasks.
@@
To address these limitations, extensive research has focused on two aspects, corresponding to the two stages in the mmWave radar signal processing pipeline: First, improving pre-processing methods to enhance the angular resolution of the point cloud, reduce false detections, and improve data quality [5, 6, 7]. Second, post-processing enhances the geometric and detailed information of the point cloud while reducing the workload on the detector. Such methods still perform worse than LiDAR point clouds in terms of accuracy and density, limiting their suitability for backend tasks in various robotic systems. Therefore, the enhancement of mmWave radar remains an open problem.
@@
Since the requirement of generating high-quality mmWave radar point clouds from raw information is reducing noise and filling sparse areas, the mmWave radar enhancement task can be regarded as a super-resolution task. Thus, the application of diffusion models to mmWave radar is quite natural. For example, Zhang et al. [7] apply RAH as the 2D representation of mmWave radar and successfully generate BEV maps of LiDAR-like quality. Luan et al. [8] then explore the BEV in mmWave radar super-resolution. Frustratingly, both RAH and BEV are not intuitive for humans due to the focus on velocity information and top view, respectively. Hence, the diffusion models pre-trained on massive natural image datasets cannot fully perform super-resolution generation on these 2D proxy representations.
@@
To overcome this problem, we propose a novel method that pioneers the application of fusing range images with image diffusion models. Unlike other representations, range images preserve more geometric and semantic information while aligning better with human perceptual patterns and natural images, which can be intuitively observed in Fig. 1. This enables the efficient transfer of knowledge from pre-trained image diffusion models to address the super-resolution task, thereby significantly enhancing the performance and efficiency of the mmWave radar enhancement.
@@
To validate the effectiveness of our approach, we conducted extensive benchmark comparisons against existing methods on both public datasets and our self-constructed datasets. The results show that our approach significantly outperforms existing methods in terms of mmWave radar point cloud quality, confirming the superior performance of our approach.
@@
    We conducted extensive benchmark comparisons with other methods on both public and self-constructed datasets, which validated the superior performance of the proposed approach.
@@ II Related Work
Traditional detection methods such as constant false alarm rate (CFAR) [9, 10, 11, 12] and multiple signal classification (MUSIC) [13] generate point clouds that are either excessively sparse or have low signal-to-noise ratios. Existing research on improving the quality of mmWave radar point clouds primarily focuses on improving pre-processing methods and post-processing methods.
@@ II-A Pre-processing Methods
Cheng et al. [5] generated ground truth Range-Doppler Maps (RDM) from LiDAR point clouds and used them to train Generative Adversarial Networks (GANs) for building detectors. While this approach improves detection performance over traditional methods, it still relies on traditional DOA estimation, leading to noisy mmWave radar point clouds. Prabhakara et al. [6] supervised a U-Net network using LiDAR point cloud labels, with low-resolution radar 2D heatmaps as input, to directly generate LiDAR-like point clouds. This approach retains real object data in the scene while eliminating part of the noise. Zhang et al. [7] supervised a diffusion model using LiDAR BEV images, which predicts LiDAR-like BEV images from paired radar RAH. This method can denoise and densify mmWave data at the same time. The two methods mentioned above use neural networks’ cross-modal learning to directly generate more accurate and dense point cloud data without DOA estimation. However, neither method can generate true 3D point clouds, which limits their applicability for higher-level tasks such as autonomous navigation in complex environments.
@@ II-B Post-processing Methods
Lu et al. [14] proposed MilliMap, which uses a GAN and takes sparse maps obtained from Bayesian grid mapping as input. This method can reconstruct dense grid maps. Geng et al. [15] first applied non-coherent integration and synthetic aperture accumulation methods to improve the density and angular resolution of radar point clouds, and then proposed the RDM network, which further filters out noise under the guidance of LiDAR point clouds. However, both methods rely on accurate ego-motion estimation from other sensors, which is difficult to obtain in extreme environments. Cai et al. [16] combined traditional signal processing with an adaptive neural network to generate high-quality indoor point clouds from mmWave reflection signals. Although this method integrates mmWave data collected from multiple devices, the resulting point clouds still lack sufficient geometric shape and detailed information. Lu et al. [17] proposed an innovative DNN-based odometry method that estimates self-motion through mmWave radar and IMU. This method jointly optimizes the transformation estimation and latent feature relationships between successive sampling intervals, effectively extracting useful motion information from complex sensor data. Zhuang et al. [18] proposed an incremental non-convex method instead of the traditional RANSAC approach for mapping with mmWave radar and inertial odometry, enabling robust and efficient self-velocity estimation from a single scan. However, both methods require the introduction of IMU sensors to achieve self-motion estimation in conjunction with mmWave radar.
In summary, existing methods either require fusion with other sensors to overcome the limitations of mmWave radar, or the generated point clouds are still not dense or accurate enough, making it challenging to support autonomous navigation or other advanced tasks in complex environments.
Figure 2: Illustration of point clouds generated from the raw ADC data of mmWave radar.
@@ III Preliminaries
III-A MmWave Radar Signal Processing
With the advancement of mmWave radar technology, it has transitioned from the conventional single-input single-output (SISO) radar to the more sophisticated multi-input multi-output (MIMO) radar. MIMO radar is outfitted with an array of multiple transmit and receive antennas, significantly improving spatial resolution, signal quality, and adaptability.
Figure 3: Diagram of the architecture of our proposed approach. (a) In training, the ground truth LiDAR point cloud and the raw mmWave radar point cloud are respectively projected into the range image $x_{0}$ and the multi-channel range image $c$. Then $x_{0}$ are corrupted to $x_{T}$ through the forward process, Finally, the neural network is trained to estimate the ground truth range image, conditioned on the paired radar multi-channel range images. (b) In inference, the neural network directly predicts $\hat{x}_{0}$ from pure Gaussian noise conditioned on the paired radar multi-channel range images. Subsequently, the predicted range image $\hat{x}_{0}$ is back-projected to obtain the final high-quality mmWave radar point cloud.
A frame of transmitted signals in mmWave radars comprises $N_{chirp}$ linear frequency-modulated chirps, with frequencies ranging from $f_{min}$ to $f_{max}$. Each chirp lasts persists for a duration of $T_{c}$ and possesses a bandwidth of $B=f_{max}-f_{min}$. The reflected signal and the transmitted signal are subsequently mixed via a mixer to produce an intermediate frequency (IF) signal, which is then sampled by a sampler to generate a 4D data cube. The four dimensions of this data cube correspond to the sampling of distinct transmitting antennas, receiving antennas, chirps, and time intervals.
The two-dimensional data from a single transmitting antenna and a single receiving antenna within the radar data cube is processed through the Fast Fourier Transform (FFT) to derive the RDM. Distinct range and Doppler bins correspond to targets situated at varying distances and velocities relative to the mmWave radar. The range of objects is determined using the formula $r=\Delta fc/2S$, where $\Delta f$ denotes the frequency shift between the received and transmitted signals, representing the frequency of the IF signal, $c$ is the speed of light, and $S$ is the slope of the FMCW signal. The maximum measurable distance is given by $r_{max}=cB/2S$, and the range resolution can be calculated as $r_{res}=c/2B$.
The velocity of the target is derived from the phase shift of the IF signal and can be expressed as $v=(\Delta\phi_{v})\lambda/(4\pi T_{c})$, where $\Delta\phi_{v}$ represents the phase shift obtained from a second FFT, commonly referred to as the Doppler FFT, and $\lambda$ is the wavelength.
The RDM obtained through the aforementioned method contains considerable noise, necessitating the use of detectors such as CFAR to identify targets. The azimuth and elevation angles of each target are determined by the data corresponding to the same position in the RDM, which is derived from different transmitting and receiving antennas. Due to the minimal difference in distance between the target and the various antennas, the frequency shift of the IF signal is small, but the phase shift remains substantial. Consequently, the azimuth and elevation angles of the target can be calculated using the formula $\theta=\arcsin(\lambda(\Delta\phi_{a}))/2\pi d)$, where $d$ represents the distance between the RX antennas, typically half the wavelength $\lambda$, and $\Delta\phi_{a}$ denotes the phase shift obtained from a third FFT, commonly referred to as the Angle FFT. Importantly, from this angle calculation formula, the angular resolution can be expressed as $\theta_{res}=\lambda/(N_{T}N_{R}d\cos(\theta))$, where $N_{T}$ and $N_{R}$ are the numbers of TX and RX antennas, respectively. This illustrates that the primary method for enhancing the angular resolution of an FMCW radar is by increasing the number of antennas.
The mmWave radar coordinates $(x,y,z,v)$ are derived from the range, velocity, azimuth angle, and elevation angle corresponding to the detected unit, as outlined above. The entire signal processing flow of the mmWave radar is depicted in Fig. 2.
III-B Diffusion Models
@@
[…]
@@ IV Methods
This section specifically details the proposed method. In recent years, image diffusion models, trained on massive image pairs, have significantly advanced the field of image super-resolution, inspiring us to leverage research from this domain to tackle point cloud super-resolution tasks. Consequently, we integrate range images with diffusion models. Specifically, the range image projected from the LiDAR point cloud is used to supervise the training of the diffusion model, after which high-quality range images are restored from noise, conditioned by multi-channel range data from mmWave radar. Ultimately, the range image is back-projected to generate high-quality mmWave radar point clouds. The system architecture is illustrated in Fig. 3.
@@ IV-A Range Image Construction
[…]
Figure 5: Qualitative comparisons of single-frame 3D point clouds conducted on the ColoRadar dataset. The LiDAR ground truth and the 3D mmWave radar point clouds generated by each method in the Hallway, Edgar Mine, Aspen Lab, and Outdoor scenes are shown above. A multi-color rendering scheme is employed for visualizing the point clouds, with red hues representing higher elevations. To enhance clarity, the visualization size of the point clouds from the RPDNet [5] and OS-CFAR [12] methods is scaled by a factor of two. Among the various CFAR methods [9, 10, 11, 12], only the OS-CFAR [12], exhibiting superior performance, is visualized.
@@ IV-B Data Representation
Furthermore, given the inherent penetration ability of mmWave radar, directly converting the raw mmWave radar point cloud data into range images results in multiple points being projected into the conical region corresponding to each pixel. This phenomenon causes point cloud data at greater radial distances to become occluded, leading to data loss and increased sparsity of the mmWave radar data. To mitigate this, we perform slicing operations along the range dimension of the mmWave radar point cloud and individually map the radar data from each slice, thus generating multi-channel range images. The process of constructing multi-channel range images is illustrated in Fig. 4.
@@ IV-C Diffusion-Based Range Image Prediction
As elaborated in Section III-B, the principal objective of the diffusion model is to learn the reverse process of the forward diffusion mechanism, which incrementally introduces noise to data samples according to the predefined noise schedule described in Eq. 2, where $s(t)=1$ and $\sigma(t)=t$. Given a set of LiDAR range images $x_{0}$ and multi-channel radar range images $c$ that are aligned both spatially and temporally, Gaussian noise is introduced and propagated to $x_{t}$. We follow the parameterization approach for the reverse process proposed by Karras et al. [19] to model the original data sample $D_{\theta}(x_{t},t,c)$ conditioned on $c$. During the inference stage, we perform deterministic sampling using the probabilistic flow (PF) ODE defined by Eq. 1 to accelerate the inference process. The Heun method is employed for iterative solving. The PF ODE is defined as follows:
@@ IV-C1 Design of the Network Architecture
13
[…]
  SO-CFAR [9] 12.841 2.671 15.5 11.070 2.705 14.9 17.160 6.154 13.3 21.547 10.80 11.5
CA-CFAR [10] 9.417 1.137 18.8 12.715 4.502 11.4 8.725 1.827 14.3 10.044 1.580 16.0
GO-CFAR [11] 8.356 1.115 19.5 11.507 4.194 11.6 8.097 2.271 14.2 9.946 1.295 15.6
OS-CFAR [12] 8.326 0.976 20.7 8.209 3.844 12.2 5.652 0.887 20.8 8.987 2.421 19.0
RPDNet [5] 2.754 0.339 34.4 2.434 0.760 29.4 2.833 0.717 31.5 8.183 2.203 24.3
  Ours 2.247 0.275 44.0 0.413 0.041 78.7 1.024 0.113 62.5 8.196 2.004 25.2
TABLE II: Quantitative results on the self-constructed dataset.
The bold denotes the best performance.
@@
7
[…]
  SO-CFAR [9] 15.606 2.267 5.7 18.368 2.733 4.7
CA-CFAR [10] 11.900 4.144 3.2 11.738 4.459 4.1
GO-CFAR [11] 6.027 2.923 0.6 13.976 10.356 0.9
OS-CFAR [12] 15.928 1.736 6.4 12.562 1.361 8.6
RPDNet [5] 10.278 1.392 6.6 6.844 1.705 14.5
  Ours 4.743 1.411 29.4 2.682 0.196 32.7

### v2

@@ Abstract
Millimeter-wave (mmWave) radar has attracted significant attention in robotics and autonomous driving. However, despite the perception stability in harsh environments, the point cloud generated by mmWave radar is relatively sparse while containing significant noise, which limits its further development. Traditional mmWave radar enhancement approaches often struggle to leverage the effectiveness of diffusion models in super-resolution, largely due to the unnatural range-azimuth heatmap (RAH) or bird’s eye view (BEV) representation. To overcome this limitation, we propose a novel method that pioneers the application of fusing range images with image diffusion models, achieving accurate and dense mmWave radar point clouds that are similar to LiDAR. Benefitting from the projection that aligns with human observation, the range image representation of mmWave radar is close to natural images, allowing the knowledge from pre-trained image diffusion models to be effectively transferred, significantly improving the overall performance. Extensive evaluations on both public datasets and self-constructed datasets demonstrate that our approach provides substantial improvements, establishing a new state-of-the-art performance in generating truly three-dimensional LiDAR-like point clouds via mmWave radar. Code will be released after publication.
@@ I Introduction
Figure 1: The comparative demonstration includes LiDAR range images and point clouds, as well as the range images and high-quality millimeter-wave radar point clouds obtained through the proposed method. The red and green dashed boxes highlight the flowerbed and bench in the current scene, respectively. Humans can easily identify object categories and geometric structures through range images, demonstrating that range images preserve rich semantic and geometric information. As a proxy representation of point clouds, it enables the knowledge of pre-trained image diffusion models to be fully leveraged in point cloud super-resolution tasks.
@@
To address these limitations, extensive research has focused on two aspects, corresponding to the two stages in the mmWave radar signal processing pipeline: First, improving pre-processing methods to enhance the angular resolution of the point cloud, reduce false detections, and improve data quality [5, 6, 7]. Second, post-processing enhances the geometric and detailed information of the point cloud while reducing the workload on the detector. The point clouds obtained by existing methods still have poor accuracy and density, limiting their suitability for backend tasks in various robotic systems. Therefore, the enhancement of mmWave radar remains an open problem.
@@
Since the requirement of generating high-quality mmWave radar point clouds from raw information is to reduce noise and fill sparse areas, the mmWave radar enhancement task can be regarded as a super-resolution task. Thus, the application of diffusion models to mmWave radar is quite natural.
At present, several related studies [7, 8, 9] have explored the application of diffusion models to the generation of mmWave radar point clouds. Nearly all of these approaches employ mmWave radar data, structured in a particular format, as a conditioning input to guide the diffusion process. For example, Zhang et al. [7] apply RAH as the 2D representation of mmWave radar and successfully generate BEV maps of LiDAR-like quality. Some studies [8, 9] have also explored converting raw mmWave radar point clouds into BEV to steer diffusion models in the generation of dense point clouds. These methods, including RAH and BEV, emphasize velocity information and top-down perspectives, which, unlike natural images, lack abundant pixel-level structural details and semantic context. However, diffusion models customarily depend on such information to achieve effective representation learning. Consequently, there is an urgent need for a more suitable representation of mmWave radar data to enable diffusion models to adapt effectively to these inputs.
@@
To overcome this problem, we propose a novel approach that, for the first time, integrates range images with diffusion models to generate mmWave radar point clouds. Unlike other representations, range images preserve more geometric and semantic information while aligning better with human perceptual patterns and natural images, which can be intuitively observed in Fig. 1. This enables the diffusion model to seamlessly adapt to mmWave radar data and perform representation learning with greater efficiency, thereby markedly enhancing the performance of mmWave radar reconstruction.
@@
To validate the performance of our approach, we conducted extensive benchmark comparisons with existing methods on public datasets and seamlessly transferred the models trained on these datasets to our self-constructed datasets. The results indicate that our method markedly surpasses existing techniques in mmWave radar point cloud quality and exhibits some generalization capability, demonstrating its superior performance.
@@
    We conducted extensive benchmark, generalization and ablation experiments on both public and self-constructed datasets, which validated the superior performance of the proposed approach.
@@ II Related Work
Traditional detection methods, such as constant false alarm rate (CFAR) [10, 11, 12, 13] and multiple signal classification (MUSIC) [14], generate point clouds that are either excessively sparse or have low signal-to-noise ratios. Existing research on improving the quality of mmWave radar point clouds primarily focuses on improving pre-processing methods and post-processing methods.
@@ II-A Pre-processing Methods
Cheng et al. [5] generated ground truth Range-Doppler Maps (RDM) from LiDAR point clouds and used them to train Generative Adversarial Networks (GANs) for building detectors. While this approach improves detection performance over traditional methods, it still relies on traditional DOA estimation, leading to noisy mmWave radar point clouds. Prabhakara et al. [6] supervised a U-Net network using LiDAR point cloud labels, with low-resolution radar 2D heatmaps as input, to directly generate LiDAR-like point clouds. This approach retains real object data in the scene while eliminating part of the noise. Han et al. [15] introduced the DenserRadar network, which directly processes the massive raw mmWave radar 4D cube using a three-dimensional U-Net architecture. However, its computational overhead is enormous. Zhang et al. [7] supervised a diffusion model using LiDAR BEV images, which predicts LiDAR-like BEV images from paired radar RAH. The methods mentioned above use neural networks’ cross-modal learning to directly generate more accurate and dense point cloud data without DOA estimation. In addition to approaches founded on LiDAR supervision, there are also methods that employ visual-inertial systems to produce ground truth labels for the purpose of cross-modal supervision. For example, Fan et al. [16] generated labels using a dynamic 3D reconstruction algorithm, thereby replacing the LiDAR-based labels in RPDNet and attaining performance on par with LiDAR-based supervision. However, these methods either fall short of producing authentic 3D point clouds or yield point clouds that remain comparatively sparse, thereby constraining their suitability for more advanced applications, such as autonomous navigation within complex environments.
@@ II-B Post-processing Methods
Lu et al. [17] proposed MilliMap, which uses a GAN and takes sparse maps obtained from Bayesian grid mapping as input. This method can reconstruct dense grid maps. Geng et al. [18] first applied non-coherent integration and synthetic aperture accumulation methods to improve the density and angular resolution of radar point clouds, and then proposed the RDM network, which further filters out noise under the guidance of LiDAR point clouds. However, both methods rely on accurate ego-motion estimation from other sensors, which is difficult to obtain in extreme environments. Cai et al. [19] combined traditional signal processing with an adaptive neural network to generate high-quality indoor point clouds from mmWave reflection signals. Although this method integrates mmWave data collected from multiple devices, the resulting point clouds still lack sufficient geometric shape and detailed information. Wu et al. [9] projected the raw mmWave radar point cloud into a BEV representation and extracted features as conditioning inputs to guide the diffusion model in reconstructing the point cloud. However, this data structure lacks abundant structural and semantic information, preventing the diffusion model from performing efficient representation learning.
In summary, existing methods either require fusion with other sensors to overcome the limitations of mmWave radar, or produce point clouds that are still not dense or accurate enough, making it challenging to support autonomous navigation or other advanced tasks in complex environments
@@ III Preliminaries
Figure 2: Diagram of the architecture of our proposed approach. (a) In training, the ground truth LiDAR point cloud and the raw mmWave radar point cloud are respectively projected into the range image $x_{0}$ and the multi-channel range image $c$. Then $x_{0}$ is corrupted to $x_{T}$ through the forward process. Finally, the neural network is trained to estimate the ground truth range image, conditioned on the paired radar multi-channel range images. (b) In inference, the neural network directly predicts $\hat{x}_{0}$ from pure Gaussian noise conditioned on the paired radar multi-channel range images. Subsequently, the predicted range image $\hat{x}_{0}$ is back-projected to obtain the final high-quality mmWave radar point cloud.
III-A Diffusion Models
@@
where $q(\cdot|\cdot)$ represents the conditional probability, $\mathcal{N}(x;\mu,\Sigma)$ denotes the probability density function of $\mathcal{N}(\mu,\Sigma)$ evaluated at $x$, $t\in\{1,\dots,T\}$, where $T$ is the total number of steps in the diffusion chain, $s(t)=\exp\left(\int_{0}^{t}f(\xi)\,d\xi\right)$, and $\sigma(t)=\sqrt{\int_{0}^{t}\frac{g(\xi)^{2}}{s(\xi)^{2}}\,d\xi}$. Gaussian noise can be added to $x_{0}$ to yield any $x_{t}$ in a single step. The forward process models a fixed Markov chain, and the noise depends on a variance schedule $s^{2}(t)\sigma^{2}(t)$.
@@ IV Methods
This section specifically details the proposed method. In recent years, image diffusion models, trained on massive image pairs, have significantly advanced the field of image super-resolution, inspiring us to leverage research from this domain to tackle point cloud super-resolution tasks. Consequently, we integrate range images with diffusion models. Specifically, the range image projected from the LiDAR point cloud is used to supervise the training of the diffusion model, after which high-quality range images are restored from noise, conditioned by range data from mmWave radar. Ultimately, the range image is back-projected to generate high-quality mmWave radar point clouds. The system architecture is illustrated in Fig. 2.
@@ IV-A Range Image Construction
where $r=\|\mathbf{p_{i}}\|_{2}$ denotes the range of each point, $\theta_{\text{min }}$ and $\phi_{\text{min }}$ represent the minimum values of the azimuth and elevation angles, $r_{\theta}$ and $r_{\phi}$ are the ranges of the azimuth and elevation angles, $l_{w}$ and $l_{h}$ are the width and height of the range image, and $\lfloor\cdot\rfloor$ denotes the floor function.
Figure 4: Qualitative comparisons of single-frame 3D point clouds conducted on the ColoRadar dataset. The LiDAR ground truth and the 3D mmWave radar point clouds generated by each method in the Hallway, Edgar Mine, Aspen Lab, and Outdoor scenes are shown above. A multi-color rendering scheme is employed for visualizing the point clouds, with red hues representing higher elevations.
@@ IV-B Data Representation
Furthermore, given the inherent penetration ability of mmWave radar, directly converting the raw mmWave radar point cloud data into range images results in multiple points being projected into the conical region corresponding to each pixel. This phenomenon causes point cloud data at greater radial distances to become occluded, leading to data loss and increased sparsity of the mmWave radar data. To mitigate this, we perform slicing operations along the range dimension of the mmWave radar point cloud and individually map the radar data from each slice, thus generating multi-channel range images as the conditions of the diffusion model. The necessity of multiple channels is also validated in Section V-B2. The process of constructing multi-channel range images is illustrated in Fig. 3.
@@ IV-C Diffusion-Based Range Image Prediction
As elaborated in Section III-A, the principal objective of the diffusion model is to learn the reverse process of the forward diffusion mechanism, which incrementally introduces noise to data samples according to the predefined noise schedule described in Eq. 2, where $s(t)=1$ and $\sigma(t)=t$. Given a set of LiDAR range images $x_{0}$ and multi-channel radar range images $c$ that are aligned both spatially and temporally, Gaussian noise is introduced and propagated to $x_{t}$. We follow the parameterization approach for the reverse process proposed by Karras et al. [20] to model the original data sample $D_{\theta}(x_{t},t,c)$ conditioned on $c$. During the inference stage, we perform deterministic sampling using the probabilistic flow (PF) ODE defined by Eq. 1 to accelerate the inference process. The Heun method is employed for iterative solving. The PF ODE is defined as follows:
@@ IV-C1 Design of the Network Architecture
CD(m) $\downarrow$ MHD(m) $\downarrow$ F-Score($\%$) $\uparrow$ CD(m) $\downarrow$ MHD(m) $\downarrow$ F-Score($\%$) $\uparrow$ CD(m) $\downarrow$ MHD(m) $\downarrow$ F-Score($\%$) $\uparrow$ CD(m) $\downarrow$ MHD(m) $\downarrow$ F-Score($\%$) $\uparrow$
  OS-CFAR [13] 5.741 0.902 20.8 8.327 0.963 20.8 4.79 1.064 25.6 9.20 2.459 18.9
MUSIC [14] 11.782 0.722 4.574 18.277 0.694 4.006 17.605 1.054 1.6 13.653 2.038 4.109
RPDNet [5] 2.329 0.545 32.8 3.044 0.464 33.7 2.927 0.858 30.7 8.183 2.203 24.3
DenserRadar [15] 8.345 4.081 14.8 3.603 0.653 30.7 2.315 0.742 37.4 20.2 11.5 15.7
  Ours 2.259 0.522 48.6 2.754 0.644 36.3 2.042 0.676 48.8 6.615 1.948 26
TABLE II: Generalization testing of the Proposed Method.
@@
CD(m) $\downarrow$ MHD(m) $\downarrow$ F-Score($\%$) $\uparrow$ CD(m) $\downarrow$ MHD(m) $\downarrow$ F-Score($\%$) $\uparrow$
  OS-CFAR [13] 29.712 3.08 2.1 17.294 3.713 2.5
Ours 7.615 1.058 11.4 8.835 2.584 7.6